OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=3/2x y=-1/2x+4

OpenStudy (er.mohd.amir):

multiply second equation by 3 and add

OpenStudy (er.mohd.amir):

3y =-3/2 x +12 and y=3/2 x

OpenStudy (anonymous):

i cant type fractions :P